# Efficient Model-Stealing Attacks Against Inductive Graph Neural Networks

### v2

@@ 6.2 Evaluation of Type I Attacks
Accuracy Fidelity Accuracy Fidelity Accuracy Fidelity
Shen et al. 0.731$\pm$0.000 0.822$\pm$0.000 0.687$\pm$0.003 0.761$\pm$0.004 0.732$\pm$0.001 0.817$\pm$0.000
Prediction ours 0.747$\pm$0.001 0.813$\pm$0.001 0.746$\pm$0.005 0.815$\pm$0.003 0.741$\pm$0.005 0.804$\pm$0.005
Shen et al. 0.681$\pm$0.001 0.752$\pm$0.000 0.680$\pm$0.004 0.752$\pm$0.005 0.677$\pm$0.000 0.751$\pm$0.001
Projection ours 0.706$\pm$0.003 0.767$\pm$0.003 0.697$\pm$0.007 0.753$\pm$0.005 0.693$\pm$0.006 0.756$\pm$0.011
Shen et al. 0.720$\pm$0.006 0.796$\pm$0.000 0.682$\pm$0.003 0.751$\pm$0.001 0.741$\pm$0.002 0.818$\pm$0.003
DBLP Embedding ours 0.742$\pm$0.002 0.785$\pm$0.002 0.749$\pm$0.005 0.791$\pm$0.002 0.704$\pm$0.011 0.744$\pm$0.009
Shen et al. 0.786$\pm$0.001 0.831$\pm$0.000 0.731$\pm$0.005 0.785$\pm$0.003 0.785$\pm$0.001 0.820$\pm$0.000
Prediction ours 0.853$\pm$0.001 0.889$\pm$0.001 0.834$\pm$0.004 0.887$\pm$0.005 0.848$\pm$0.001 0.884$\pm$0.001
Shen et al. 0.814$\pm$0.002 0.874$\pm$0.000 0.798$\pm$0.001 0.871$\pm$0.001 0.822$\pm$0.000 0.878$\pm$0.000
Projection ours 0.846$\pm$0.000 0.884$\pm$0.002 0.830$\pm$0.001 0.883$\pm$0.002 0.829$\pm$0.004 0.866$\pm$0.005
Shen et al. 0.831$\pm$0.002 0.887$\pm$0.000 0.821$\pm$0.002 0.889$\pm$0.006 0.818$\pm$0.000 0.864$\pm$0.001
Pubmed Embedding ours 0.855$\pm$0.001 0.879$\pm$0.001 0.849$\pm$0.002 0.883$\pm$0.002 0.850$\pm$0.002 0.868$\pm$0.003
Shen et al. 0.797$\pm$0.001 0.854$\pm$0.000 0.764$\pm$0.001 0.815$\pm$0.001 0.797$\pm$0.001 0.853$\pm$0.001
Prediction ours 0.813$\pm$0.002 0.854$\pm$0.004 0.840$\pm$0.005 0.887$\pm$0.004 0.813$\pm$0.003 0.856$\pm$0.005
Shen et al. 0.660$\pm$0.006 0.696$\pm$0.000 0.709$\pm$0.005 0.754$\pm$0.003 0.651$\pm$0.010 0.680$\pm$0.008
Projection ours 0.727$\pm$0.005 0.747$\pm$0.002 0.757$\pm$0.002 0.784$\pm$0.001 0.667$\pm$0.048 0.691$\pm$0.051
Shen et al. 0.780$\pm$0.002 0.817$\pm$0.000 0.770$\pm$0.002 0.796$\pm$0.004 0.787$\pm$0.003 0.818$\pm$0.001
Citeseer Embedding ours 0.802$\pm$0.006 0.825$\pm$0.006 0.838$\pm$0.004 0.851$\pm$0.009 0.791$\pm$0.005 0.800$\pm$0.005
Shen et al. 0.881$\pm$0.002 0.828$\pm$0.000 0.876$\pm$0.002 0.839$\pm$0.005 0.642$\pm$0.006 0.586$\pm$0.007
Prediction ours 0.911$\pm$0.005 0.833$\pm$0.002 0.874$\pm$0.024 0.813$\pm$0.034 0.829$\pm$0.008 0.775$\pm$0.014
Shen et al. 0.532$\pm$0.004 0.479$\pm$0.000 0.480$\pm$0.116 0.455$\pm$0.122 0.500$\pm$0.009 0.461$\pm$0.005
Projection ours 0.661$\pm$0.028 0.602$\pm$0.029 0.462$\pm$0.028 0.418$\pm$0.025 0.473$\pm$0.048 0.424$\pm$0.047
Shen et al. 0.744$\pm$0.005 0.683$\pm$0.000 0.863$\pm$0.002 0.842$\pm$0.001 0.525$\pm$0.006 0.473$\pm$0.004
Amazon Embedding ours 0.914$\pm$0.003 0.838$\pm$0.006 0.901$\pm$0.004 0.851$\pm$0.006 0.893$\pm$0.008 0.811$\pm$0.003
Shen et al. 0.942$\pm$0.000 0.958$\pm$0.000 0.889$\pm$0.008 0.921$\pm$0.008 0.944$\pm$0.000 0.961$\pm$0.000
Prediction ours 0.946$\pm$0.002 0.952$\pm$0.001 0.888$\pm$0.002 0.917$\pm$0.002 0.944$\pm$0.000 0.953$\pm$0.001
Shen et al. 0.848$\pm$0.002 0.856$\pm$0.000 0.850$\pm$0.018 0.866$\pm$0.018 0.813$\pm$0.001 0.824$\pm$0.001
Projection ours 0.933$\pm$0.001 0.940$\pm$0.000 0.734$\pm$0.008 0.750$\pm$0.009 0.873$\pm$0.018 0.884$\pm$0.015
Shen et al. 0.946$\pm$0.001 0.959$\pm$0.000 0.923$\pm$0.001 0.944$\pm$0.001 0.922$\pm$0.002 0.936$\pm$0.004
Coauthor Embedding ours 0.941$\pm$0.001 0.946$\pm$0.001 0.898$\pm$0.005 0.925$\pm$0.005 0.936$\pm$0.004 0.941$\pm$0.003
Shen et al. 0.905$\pm$0.000 0.945$\pm$0.000 0.892$\pm$0.000 0.959$\pm$0.000 0.897$\pm$0.000 0.943$\pm$0.000
Prediction ours 0.907$\pm$0.005 0.913$\pm$0.003 0.900$\pm$0.002 0.926$\pm$0.002 0.892$\pm$0.011 0.911$\pm$0.016
Shen et al. 0.835$\pm$0.000 0.858$\pm$0.000 0.872$\pm$0.000 0.920$\pm$0.000 0.876$\pm$0.000 0.904$\pm$0.000
Projection ours 0.864$\pm$0.009 0.879$\pm$0.005 0.855$\pm$0.008 0.884$\pm$0.007 0.890$\pm$0.009 0.895$\pm$0.014
Shen et al. 0.906$\pm$0.000 0.935$\pm$0.000 0.830$\pm$0.000 0.871$\pm$0.000 0.861$\pm$0.000 0.882$\pm$0.000
ACM Embedding ours 0.885$\pm$0.007 0.902$\pm$0.008 0.882$\pm$0.009 0.914$\pm$0.008 0.868$\pm$0.006 0.887$\pm$0.001
@@
The overview of the accuracy of the original node classification tasks for the target models was shown in Table 2. All analyzed GNN models perform well across all datasets, underscoring the effectiveness of jointly considering node features and graph structure for classification. Subsequently, the accuracy and fidelity results for Type I attacks are presented in Table 3. We focus on the results when the adversary targets the GIN model with the stealing attack. Similar performance patterns with GAT and Sage as target models are detailed in Appendix C due to space constraints.
@@
Accuracy. In our analysis of the target GIN model, a noteworthy surge in accuracy is observed across all six datasets when utilizing various surrogate models for the prediction task. Particularly remarkable is the ACM dataset, where accuracy registers a substantial increase of approximately 15 %pt. across all cases. This trend extends to the projection task, showcasing enhanced accuracy for most datasets, including DBLP, Pubmed, Citeseer, and Coauthor. Furthermore, our method exhibits a significant improvement in embedding accuracy across all datasets and surrogate models. Remarkably, our approach outperforms the baseline method in terms of accuracy for all tasks, datasets, and surrogate models, with the most pronounced benefits evident in the DBLP dataset.
@@ 6.3 Evaluation of Type II Attacks
Accuracy. Under the GIN target model, our approach consistently enhances prediction accuracy across the majority of datasets and surrogate models. Noteworthy is the 18.7 %pt. increase observed with the SAGE surrogate model on the Amazon dataset. This trend persists in the context of projection accuracy, where our method outperforms competing configurations. Similarly, substantial improvements in embedding accuracy are evident, particularly on the Amazon dataset with the SAGE surrogate model, showcasing a remarkable 35.8 %pt. lead over the baseline.
@@ 6.4 Query Budget
We explore the efficacy of our attack under varying query budgets, represented by different sizes of the query graph $\mathbf{G}_{Q}$. We compare the query efficiency of our approach to that of [30]. Due to constraints on available space, we present results exclusively for the Citeseer dataset; however, similar trends are observed across other datasets. We depict accuracy and fidelity outcomes for both Type I and Type II attacks for the embedding (see Figure 3) task for 3 surrogate models: GIN, GAT, SAGE. Corresponding plots for prediction and projection (see Figure 5, and Figure 6, respectively) can be found in Appendix E. We compare our results to the accuracy and fidelity scores achieved with the best-performing surrogate model type trained on 30% of dataset nodes using the method introduced by [30]. Our method achieves the same or higher level of accuracy and fidelity for prediction and embedding responses using only 50% of the queries required in the baseline method. For the projection task, we can make a similar observation for the GAT surrogate model. Learning from the t-SNE projection requires using a number of queries closer to 30% to match the performance best surrogate model of the baseline when GIN or SAGE is selected as the surrogate model.
@@ 7 Discussion
Limitations. Firstly, the proposed model stealing attack is limited to settings where the target model provides node-level results. The scenarios in which the target model accepts arbitrary graphs as input and returns graph-level results were not the subject of consideration. Secondly, the learning process of the missing graph structure was based on the approach presented in the paper [30]. Developing a novel technique for reconstructing graph structure, specifically designed for training the surrogate model with contrastive loss, constitutes a direction for further research.
Defense. Several countermeasures against model stealing attacks have been proposed in the literature [4, 6, 31]. The applied formulation of the defense mechanism follows the approach by Shen and co-authors [30]. To address diverse query responses, we inject random Gaussian noise into node embeddings and t-SNE projections produced by target models, employing GIN as the surrogate model to assess the countermeasure’s efficacy. The surrogate model’s accuracy, serving as a metric for attack performance within the defined threat model, is evaluated on the ACM dataset (refer to Figure 4 for the embedding response and Appendix F for prediction and projection responses). Our findings indicate a marginal impact of random Gaussian noise on the surrogate model’s accuracy, similar to the results obtained by [30]. Even with a high standard deviation of the noise added, the model stealing attack still results in a useful surrogate model. Moreover, injecting random noise into the target’s model responses results in a lowered utility of the model for legitimate users of the model API, as outlined in [5]. Developing robust defense mechanisms continues to be a focus for future research.

### v4

@@ 6.1 Evaluation of Type I Attacks
Accuracy Fidelity F1 score Accuracy Fidelity F1 score Accuracy Fidelity F1 score
Shen et al. 0.731$\pm$0.000 0.822$\pm$0.000 0.565$\pm$0.004 0.687$\pm$0.003 0.761$\pm$0.004 0.453$\pm$0.003 0.732$\pm$0.001 0.817$\pm$0.000 0.439$\pm$0.003
Prediction ours 0.747$\pm$0.001 0.813$\pm$0.001 0.628$\pm$0.004 0.746$\pm$0.005 0.815$\pm$0.003 0.682$\pm$0.005 0.741$\pm$0.005 0.804$\pm$0.005 0.648$\pm$0.003
Shen et al. 0.681$\pm$0.001 0.752$\pm$0.000 0.366$\pm$0.001 0.680$\pm$0.004 0.752$\pm$0.005 0.518$\pm$0.003 0.677$\pm$0.000 0.751$\pm$0.001 0.362$\pm$0.003
Projection ours 0.706$\pm$0.003 0.767$\pm$0.003 0.574$\pm$0.001 0.697$\pm$0.007 0.753$\pm$0.005 0.505$\pm$0.001 0.693$\pm$0.006 0.756$\pm$0.011 0.579$\pm$0.002
Shen et al. 0.720$\pm$0.006 0.796$\pm$0.000 0.540$\pm$0.005 0.682$\pm$0.003 0.751$\pm$0.001 0.540$\pm$0.004 0.741$\pm$0.002 0.818$\pm$0.003 0.605$\pm$0.005
DBLP Embedding ours 0.742$\pm$0.002 0.785$\pm$0.002 0.650$\pm$0.005 0.749$\pm$0.005 0.791$\pm$0.002 0.677$\pm$0.004 0.704$\pm$0.011 0.744$\pm$0.009 0.613$\pm$0.001
Shen et al. 0.786$\pm$0.001 0.831$\pm$0.000 0.738$\pm$0.001 0.731$\pm$0.005 0.785$\pm$0.003 0.704$\pm$0.001 0.785$\pm$0.001 0.820$\pm$0.000 0.691$\pm$0.003
Prediction ours 0.853$\pm$0.001 0.889$\pm$0.001 0.847$\pm$0.004 0.834$\pm$0.004 0.887$\pm$0.005 0.822$\pm$0.003 0.848$\pm$0.001 0.884$\pm$0.001 0.846$\pm$0.004
Shen et al. 0.814$\pm$0.002 0.874$\pm$0.000 0.716$\pm$0.001 0.798$\pm$0.001 0.871$\pm$0.001 0.801$\pm$0.001 0.822$\pm$0.000 0.878$\pm$0.000 0.687$\pm$0.003
Projection ours 0.846$\pm$0.000 0.884$\pm$0.002 0.836$\pm$0.003 0.830$\pm$0.001 0.883$\pm$0.002 0.805$\pm$0.009 0.829$\pm$0.004 0.866$\pm$0.005 0.821$\pm$0.002
Shen et al. 0.831$\pm$0.002 0.887$\pm$0.000 0.774$\pm$0.002 0.821$\pm$0.002 0.889$\pm$0.006 0.819$\pm$0.003 0.818$\pm$0.000 0.864$\pm$0.001 0.676$\pm$0.001
Pubmed Embedding ours 0.855$\pm$0.001 0.879$\pm$0.001 0.857$\pm$0.001 0.849$\pm$0.002 0.883$\pm$0.002 0.839$\pm$0.004 0.850$\pm$0.002 0.868$\pm$0.003 0.842$\pm$0.001
Shen et al. 0.797$\pm$0.001 0.854$\pm$0.000 0.810$\pm$0.004 0.764$\pm$0.001 0.815$\pm$0.001 0.820$\pm$0.009 0.797$\pm$0.001 0.853$\pm$0.001 0.798$\pm$0.001
Prediction ours 0.813$\pm$0.002 0.854$\pm$0.004 0.822$\pm$0.001 0.840$\pm$0.005 0.887$\pm$0.004 0.836$\pm$0.001 0.813$\pm$0.003 0.856$\pm$0.005 0.813$\pm$0.003
Shen et al. 0.660$\pm$0.006 0.696$\pm$0.000 0.670$\pm$0.003 0.709$\pm$0.005 0.754$\pm$0.003 0.667$\pm$0.001 0.651$\pm$0.010 0.680$\pm$0.008 0.675$\pm$0.004
Projection ours 0.727$\pm$0.005 0.747$\pm$0.002 0.714$\pm$0.009 0.757$\pm$0.002 0.784$\pm$0.001 0.731$\pm$0.001 0.667$\pm$0.048 0.691$\pm$0.051 0.583$\pm$0.005
Shen et al. 0.780$\pm$0.002 0.817$\pm$0.000 0.791$\pm$0.009 0.770$\pm$0.002 0.796$\pm$0.004 0.760$\pm$0.001 0.787$\pm$0.003 0.818$\pm$0.001 0.803$\pm$0.001
Citeseer Embedding ours 0.802$\pm$0.006 0.825$\pm$0.006 0.800$\pm$0.002 0.838$\pm$0.004 0.851$\pm$0.009 0.844$\pm$0.001 0.791$\pm$0.005 0.800$\pm$0.005 0.808$\pm$0.009
Shen et al. 0.881$\pm$0.002 0.828$\pm$0.000 0.833$\pm$0.002 0.876$\pm$0.002 0.839$\pm$0.005 0.575$\pm$0.001 0.642$\pm$0.006 0.586$\pm$0.007 0.680$\pm$0.009
Prediction ours 0.911$\pm$0.005 0.833$\pm$0.002 0.900$\pm$0.002 0.874$\pm$0.024 0.813$\pm$0.034 0.882$\pm$0.003 0.829$\pm$0.008 0.775$\pm$0.014 0.774$\pm$0.001
Shen et al. 0.532$\pm$0.004 0.479$\pm$0.000 0.654$\pm$0.009 0.480$\pm$0.116 0.455$\pm$0.122 0.301$\pm$0.005 0.500$\pm$0.009 0.461$\pm$0.005 0.688$\pm$0.001
Projection ours 0.661$\pm$0.028 0.602$\pm$0.029 0.673$\pm$0.001 0.462$\pm$0.028 0.418$\pm$0.025 0.492$\pm$0.002 0.473$\pm$0.048 0.424$\pm$0.047 0.491$\pm$0.004
Shen et al. 0.744$\pm$0.005 0.683$\pm$0.000 0.847$\pm$0.001 0.863$\pm$0.002 0.842$\pm$0.001 0.619$\pm$0.004 0.525$\pm$0.006 0.473$\pm$0.004 0.725$\pm$0.002
Amazon Embedding ours 0.914$\pm$0.003 0.838$\pm$0.006 0.883$\pm$0.009 0.901$\pm$0.004 0.851$\pm$0.006 0.880$\pm$0.005 0.893$\pm$0.008 0.811$\pm$0.003 0.835$\pm$0.009
Shen et al. 0.942$\pm$0.000 0.958$\pm$0.000 0.910$\pm$0.005 0.889$\pm$0.008 0.921$\pm$0.008 0.910$\pm$0.009 0.944$\pm$0.000 0.961$\pm$0.000 0.931$\pm$0.005
Prediction ours 0.946$\pm$0.002 0.952$\pm$0.001 0.920$\pm$0.001 0.888$\pm$0.002 0.917$\pm$0.002 0.910$\pm$0.001 0.944$\pm$0.000 0.953$\pm$0.001 0.933$\pm$0.003
Shen et al. 0.848$\pm$0.002 0.856$\pm$0.000 0.876$\pm$0.005 0.850$\pm$0.018 0.866$\pm$0.018 0.861$\pm$0.001 0.813$\pm$0.001 0.824$\pm$0.001 0.822$\pm$0.002
Projection ours 0.933$\pm$0.001 0.940$\pm$0.000 0.911$\pm$0.009 0.734$\pm$0.008 0.750$\pm$0.009 0.850$\pm$0.002 0.873$\pm$0.018 0.884$\pm$0.015 0.890$\pm$0.001
Shen et al. 0.946$\pm$0.001 0.959$\pm$0.000 0.916$\pm$0.004 0.923$\pm$0.001 0.944$\pm$0.001 0.930$\pm$0.004 0.922$\pm$0.002 0.936$\pm$0.004 0.949$\pm$0.001
Coauthor Embedding ours 0.941$\pm$0.001 0.946$\pm$0.001 0.916$\pm$0.004 0.898$\pm$0.005 0.925$\pm$0.005 0.920$\pm$0.001 0.936$\pm$0.004 0.941$\pm$0.003 0.960$\pm$0.003
Shen et al. 0.905$\pm$0.000 0.945$\pm$0.000 0.901$\pm$0.009 0.892$\pm$0.000 0.959$\pm$0.000 0.876$\pm$0.002 0.897$\pm$0.000 0.943$\pm$0.000 0.900$\pm$0.001
Prediction ours 0.907$\pm$0.005 0.913$\pm$0.003 0.904$\pm$0.002 0.900$\pm$0.002 0.926$\pm$0.002 0.894$\pm$0.001 0.892$\pm$0.011 0.911$\pm$0.016 0.910$\pm$0.002
Shen et al. 0.835$\pm$0.000 0.858$\pm$0.000 0.884$\pm$0.002 0.872$\pm$0.000 0.920$\pm$0.000 0.859$\pm$0.003 0.876$\pm$0.000 0.904$\pm$0.000 0.879$\pm$0.001
Prediction ours 0.907$\pm$0.005 0.913$\pm$0.003 0.904$\pm$0.002 0.900$\pm$0.002 0.926$\pm$0.002 0.894$\pm$0.001 0.892$\pm$0.011 0.911$\pm$0.016 0.910$\pm$0.002
Shen et al. 0.906$\pm$0.000 0.935$\pm$0.000 0.910$\pm$0.005 0.830$\pm$0.000 0.871$\pm$0.000 0.875$\pm$0.002 0.861$\pm$0.000 0.882$\pm$0.000 0.876$\pm$0.004
ACM Embedding ours 0.885$\pm$0.007 0.902$\pm$0.008 0.880$\pm$0.001 0.882$\pm$0.009 0.914$\pm$0.008 0.879$\pm$0.002 0.868$\pm$0.006 0.887$\pm$0.001 0.884$\pm$0.001
@@
The overview of the accuracy and F1 score of the original node classification tasks for the target models are shown in Table 2. All analyzed GNN models perform well across all datasets, underscoring the effectiveness of jointly considering node features and graph structure for classification. Subsequently, the accuracy and fidelity results for Type I attacks are presented in Table 3. We focus on the results when the adversary targets the GIN model with the stealing attack. Similar performance patterns with GAT and Sage as target models are detailed in Appendix C due to space constraints.
@@
Accuracy and F1 score. In our analysis of the target GIN model, a noteworthy surge in accuracy and F1 score is observed across all six datasets when utilizing various surrogate models for the prediction task. Particularly remarkable is the ACM dataset, where accuracy and F1 register a substantial increase of approximately 15 %pt. across all cases. This trend extends to the projection task, showcasing enhanced accuracy for most datasets, including DBLP, Pubmed, Citeseer, and Coauthor. Furthermore, our method exhibits a significant improvement in embedding accuracy across all datasets and surrogate models. Remarkably, our approach outperforms the baseline method for all tasks, datasets, and surrogate models, with the most pronounced benefits evident in the DBLP dataset.
@@ 6.2 Evaluation of Type II Attacks
Accuracy and F1 score. Under the GIN target model, our approach consistently enhances prediction accuracy and F1 score across the majority of datasets and surrogate models. Noteworthy, the accuracy observed with the SAGE surrogate model on the Amazon dataset increases by 18.7 %pt. This trend persists in the context of projection accuracy, where our method outperforms competing configurations. Similarly, improvements in embedding accuracy and F1 score are evident, particularly on the Amazon dataset with the SAGE surrogate model, showcasing a remarkable 35.8 %pt. accuracy lead over the baseline.
@@ 6.3 Query Budget
We explore the efficacy of our attack under varying query budgets, represented by different sizes of the query graph $\mathbf{G}_{Q}$. We compare the query efficiency of our approach with that of [33], focusing on the Citeseer dataset due to space constraints, though similar trends are observed in other datasets. Accuracy and fidelity outcomes for both Type I and Type II attacks on the embedding task (see Figure 3) are presented for three surrogate models: GIN, GAT, and SAGE. Plots for prediction and projection tasks are provided in Appendix E. We also compare our results to the accuracy and fidelity scores achieved by the best-performing surrogate model trained on 30% of dataset nodes using the method of [33]. Our method achieves the same or higher level of accuracy and fidelity for prediction and embedding responses using only 50% of the queries required in the baseline method. For the projection task, we can make a similar observation for the GAT surrogate model. Learning from the t-SNE projection requires using a number of queries closer to 30% to match the performance best surrogate model of the baseline when GIN or SAGE is selected as the surrogate model.
@@ 7 Discussion
Limitations. Firstly, the proposed model stealing attack is limited to scenarios where the target model provides node-level results. However, our method can be extended to graph-level and link prediction tasks by adjusting the augmentations (e.g., using node instead of link augmentations for link prediction). Secondly, the learning process of the missing graph structure was based on the approach presented in the paper [33]. Developing a novel technique for reconstructing graph structure, specifically designed for training the surrogate model with contrastive loss, constitutes a direction for further research.
Defense. Several active countermeasures against model stealing attacks have been proposed in the literature [5, 7, 34]. The applied formulation of the defense mechanism follows the approach by Shen and co-authors [33]. To address diverse query responses, we inject random Gaussian noise into node embeddings and t-SNE projections produced by target models, employing GIN as the surrogate model to assess the countermeasure’s efficacy. The surrogate model’s accuracy, serving as a metric for attack performance within the defined threat model, is evaluated on the ACM dataset (refer to Figure 4 for the embedding response and Appendix F for prediction and projection responses). Our findings indicate a marginal impact of random Gaussian noise on the surrogate model’s accuracy, similar to the results obtained by [33]. Even with a high standard deviation of the noise added, the model stealing attack still results in a useful surrogate model. Moreover, injecting random noise into the target’s model responses results in a lowered utility of the model for legitimate users of the model API, as outlined in [6]. Developing robust active defense mechanisms continues to be a focus for future research. However, we note the existence of several methods for protecting the Intellectual Property after the GNN model has already been stolen [29, 39, 44, 47, 45, 4].